# Identifiability of VAR(1) model in a stationary setting

### v2

@@ 3 Jacobian Structure of VAR(1) in the Stationary Setting
The proof is given in Appendix B. ∎
In Lemma 1, $P_{\left(i-1\right)n+j,\left(j-1\right)n+i}$ is the identify matrix exchanging the $i^{th}$ and $j^{th}$ columns of the $j^{th}$ and $i^{th}$ blocks, and hence $\mathbf{P}$ is just the indentity matrix exchanging columns $n(n-1)/2$ times. The following example is a visualization of the function of $\mathbf{P}$.
@@
The proof is based on the facts that $\overline{\mathbf{J}}_{G}$ has extra rows corresponding to $\lambda_{ij}$ where $(i,j)\notin\mathfrak{E}_{G}$, and that such $\lambda_{ij}$ equals zero in $\overline{\mathbf{J}}_{G}$. The complete proof is given in Appendix C. ∎
@@
The proof is given in Appendix D. ∎
@@
Since $\det\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is a polynomial, by continuity, it does not identically equal to 00. Therefore, the set of $\Lambda$ such that the determinant equals zero must have zero measure, i.e. $\mu_{G}\left(M_{G}^{0}\right)=0$. ∎
@@ 4 Maximal classes
In this section, we introduce the definition and properties of maximal classes and highlight their relationship with some parameters of our model. In the following sections, we will see that in general, different maximal classes could imply generic identifiability. Section 4.1 formally defines maximal classes for directed graphs, and proves the uniqueness of the set of maximal classes for a given graph. Additionally, we present an algorithm that identifies the set of maximal classes from any graph, and compare this concept with related notions in graph theory. In Section 4.2, we highlight the strong relationship between maximal classes and the covariance matrix $\Sigma$ (definition in (2)), based on which another algorithm is designed to reconstruct the set of maximal classes as well as a finite list of possible graphs from the support of $\Sigma$. It serves as a first step to the reconstruction of the graph corresponding to a stationary VAR(1) model, which even though is not the main subject of this paper, still could inspire future works.
@@ 4.1 Definition and properties
${\operatorname{deg}}_{+}(C)=|\left\{\left(v,v_{i}\right)\in\mathfrak{E}_{G}\mid i\in[p],v\in V\backslash\left\{v_{1},\cdots,v_{p}\right\}\right\}|.$
@@
  1. 1.
    there exists a directed path in $G$ between $i$ and $j$;
  2. 2.
    $\exists l\neq i,j\in V$ s.t. there are directed paths from $l$ to $i$ and $j$ respectively.
@@
Because the set of maximal classes is unique (Proposition 6), we can design Algorithm 1 to derive the set of maximal classes for any directed graph. This Algorithm first applies Kosaraju’s algorithm ([13]) to find all SCCs of the graph, and then construct the condensed graph, which is a directed acyclic graph where all SCCs are represented as a single node. Finally, we perform a Depth First Search (DFS) algorithm ([16]) on each of the nodes with in-degree zero in the condensed graph. See https://github.com/Bi-xuan/maximal_class for an R implementation of the algorithm.
Algorithm 1 Algorithm to find maximal classes of a directed graph
@@
11:         DFS$(G^{\prime},L[i],j)$triangleright Put the set of all reachable nodes from $i$ into the $j^{th}$ element of the list
12:         $j\leftarrow j+1$     return $j-1$ (number of maximal classes) and maxc (list of maximal classes)
@@ 4.2 Derivation of maximal classes from $\Sigma$
$M^{P}:=$$\left\{\Lambda\in M_{n}\left(\mathbb{R}\right)\mid\forall i,j\in V,\exists\mbox{ a directed path in }G\mbox{ from }i\mbox{ to }j\Rightarrow\exists K\in\mathbb{N},s.t.\lambda_{ij}^{\left[k\right]}\neq 0,\forall k>K\right\},$
@@
The proof is given in Appendix F. ∎
@@
$\superscript\mathfrak{MC}=\mathfrak{C}\backslash\mathfrak{C}^{\prime},$
@@
$\mathfrak{C}$ =$=\left\{\mathcal{C}_{i}\mid i\in[n]\right\},$
$\mathfrak{C}^{\prime}$ =$=\left\{\mathcal{C}_{i}\in\mathfrak{C}\mid i\in[n],\mbox{ and }\exists k,l\in\mathcal{C}_{i}\ s.t.\ \sigma_{kl}=0\right\}.$
@@
  1. 1.
    There exists a directed path from $i$ to $j$, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction;
  2. 2.
    There exists a directed path from $j$ to $i$, then ${\operatorname{deg}}_{+}(i)\neq 0$, which means that the source of $\mathcal{MC}_{[i]}$ is a SCC containing $i$. Hence either $j$ belongs to the SCC, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction, or $j$ does not belong to the SCC, then the in-degree of the SCC is non-zero, i.e. the SCC containing $i$ is not the source of $\mathcal{MC}_{[i]}$, which is also a contradiction.
  3. 3.
    There exists another node $c\in[n]$ s.t. there are paths from $c$ to $i$, and from $c$ to $j$ respectively, then ${\operatorname{deg}}_{+}(i)\neq 0$, which means that the source of $\mathcal{MC}_{[i]}$ is a SCC containing $i$. Hence either $c$ belongs to the SCC, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction, or $c$ does not belong to the SCC, then the in-degree of the SCC is non-zero, i.e. the SCC containing $i$ is not the source of $\mathcal{MC}_{[i]}$, which is also a contradiction.
@@
$\superscript\mathfrak{MC}\subseteq\mathfrak{C}\backslash\mathfrak{C}^{\prime}.$
Let $\superscript\mathcal{C}_{i}\in\mathfrak{C}\backslash\mathfrak{C}^{\prime}$ for certain $i\in[n]$ and assume that $\mathcal{C}_{i}\notin\mathfrak{MC}$ (proof in Appendix E). We know that $i$ must belong to one maximal class in $\mathfrak{MC}$. Denote the maximal class containing $i$ as $\mathcal{MC}_{i}$. Then by Proposition 8,
@@
$\superscript\mathfrak{MC}=\mathfrak{C}\backslash\mathfrak{C}^{\prime}.$
@@
Algorithm 2 Algorithm to find maximal classes from the support of $\Sigma$
@@
12:             $\mathfrak{MC}\leftarrow\mathfrak{MC}\backslash\mathfrak{MC}[i]$triangleright Keep only one of the repetitive elements
@@
=$\mathfrak{C}=\left\{\{1,2\},\{1,2,3\},\{2,3\}\right\},$
@@
=$\mathfrak{C}^{\prime}=\left\{\{1,2,3\}\right\},$
@@ 5.1 Maximal class characterization
$\psi_{G}\left(B\right)^{\left(\lambda_{st}\right)}=$ $\left[\begin{array}[]{cccc}\begin{array}[]{ccccc}0&\cdots&\left(\Sigma\Lambda\right)_{s1}&\cdots&0\end{array}&\cdots&\begin{array}[]{ccccc}\left(\Sigma\Lambda\right)_{s1}&\cdots&\left(\Sigma\Lambda\right)_{st}&\cdots&\left(\Sigma\Lambda\right)_{sn}\end{array}&\cdots\end{array}\right.$
@@
  • Case 1: without loss of generality, assume the path is from $s$ to $a$, then there exist directed paths: $s\rightsquigarrow t\rightsquigarrow j$, and $s\rightsquigarrow a\rightsquigarrow i$. It contradicts the assumption.
  • Case 2: assume there exist directed paths: $l\rightsquigarrow s$, and $l\rightsquigarrow a$, then there exist directed paths: $l\rightsquigarrow a\rightsquigarrow i$, and $l\rightsquigarrow s\rightarrow t\rightsquigarrow j$. It also contradicts the assumption.
Therefore, for all $a\in\left[n\right]$ and $k\in\mathbb{N}$, $\left(\Sigma\Lambda\right)_{sa}\lambda_{ai}^{\left[k\right]}\lambda_{tj}^{\left[k\right]}=0$. The second term of the sum is also 00 by similar arguments. Hence, for all $s,t\in\left[n\right]$ s.t. $\left(s,t\right)\in\mathfrak{E}_{G}$,
@@
The proof is given in Appendix G. ∎
@@
$n_{c}^{\prime}$ =|$=|\left\{\{a,b\}\mid a,b\in\left[n\right],a,b\mbox{ belong to thesame maximal class}\right\}|.$
@@
[…]
@@
[…]
@@
[…]
@@
  1. 1.
    there exist $i,j\in[n]$ s.t. $i,j$ belong to the same maximal class in $G_{k_{1}}$, but do not in $G_{k_{2}}$;
  2. 2.
    there exist $s,t\in[n]$ s.t. $s,t$ belong to the same maximal class in $G_{k_{2}}$, but do not in $G_{k_{1}}$;
@@ 5.2 The dimension of the model for graphs without multi-edges
${\operatorname{dim}}\left(\mathbf{M}_{G}\right)=\mbox{rank}\left(\mathbf{J}_{G}\right)=\min\left\{n_{r},n_{c}^{\prime}\right\}=9.$
The proof of Theorem 5 is technical, and aligns with generic settings throughout this paper. It is divided into two separate cases: $n_{r}\leq n_{c}^{\prime}$ and $n_{r}>n_{c}^{\prime}$. We only introduce the proof for the first case, since the proof of the second case resembles the first one. For a complete proof of the second case, see Appendix I.
@@
$\mathfrak{C}_{G}^{mc}:=\left\{\{k,l\}\mid k,l\in[n],(k,l),(l,k)\notin\mathfrak{E}_{G}^{\prime},\mbox{ and }k,l\mbox{ belong to the same maximal class}\right\},$
@@
where $\left\{k,l\right\}\in\mathfrak{C}_{G}^{mc}$. Note that $\mathfrak{C}_{G}^{mc}\neq\emptyset$ by Lemma 8. In fact, this submatrix is a block matrix after reordering rows and columns. Apply the the Guttman rank additivity formula (see [7]), this matrix is full rank if and only if $B_{G}$ is full rank, up to a multiplication of a diagonal matrix with non-zero diagonals. For complete proof, see Appendix H. ∎
@@
  1. 1.
    $\left(k,a\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$;
  2. 2.
    $\left(a,k\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$.
@@
is rational. From Appendix A, we know that the parametrization map
@@
Case 1: Consider $\Lambda_{0}=\left(\lambda_{ij}^{0}\right)$ that satisfies:
@@
Case 2: Let $G_{0}=\left(V,\mathfrak{E}_{G_{0}}\right)$ where
@@
Proof of case $1$ in Theorem 5.
This is a direct result of Lemmas 9 and 11. ∎

### v3

@@ 3 Jacobian Structure of VAR(1) in the Stationary Setting
The proof is given in Appendix C. ∎
In Lemma 1, $P_{\left(i-1\right)n+j,\left(j-1\right)n+i}$ is the identify matrix exchanging the $i^{th}$ and $j^{th}$ columns of the $j^{th}$ and $i^{th}$ blocks, and hence $\mathbf{P}$ is just the indentity matrix exchanging columns n(n1)/2n(n-1)/2 times. The following example is a visualization of the function of $\mathbf{P}$.
@@
The proof is based on the facts that $\overline{\mathbf{J}}_{G}$ has extra rows corresponding to $\lambda_{ij}$ where $(i,j)\notin\mathfrak{E}_{G}$, and that such $\lambda_{ij}$ equals zero in $\overline{\mathbf{J}}_{G}$. The complete proof is given in Appendix D. ∎
@@
The proof is given in Appendix E. ∎
@@
Since $\det\left(I_{n^{2}}-\Lambda\otimes\Lambda\right)$ is a polynomial, by continuity, it does not identically equal to 0. Therefore, the set of $\Lambda$ such that the determinant equals zero must have zero measure, i.e. $\mu_{G}\left(M_{G}^{0}\right)=0$. ∎
@@ 4 Maximal classes
In this section, we introduce the definition and properties of maximal classes and highlight their relationship with some parameters of our model. In the following sections, we will see that in general, different maximal classes could imply generic identifiability. Section 4.1 formally defines maximal classes for directed graphs, and proves the uniqueness of the set of maximal classes for a given graph. Additionally, we propose Algorithm 1 to identify the set of maximal classes from any graph, and Algorithm 2 to reconstruct the list of graphs associated with a given set of maximal classes. We also compare the concept of maximal classes with related notions in graph theory. In Section 4.2, we highlight the strong relationship between maximal classes and the covariance matrix of the stationary distribution $\Sigma$, based on which another algorithm (Algorithm 3) is designed to reconstruct the set of maximal classes from the support of $\Sigma$. With Algorithm 2, the list of graphs associated with $\Sigma$ can also be reconstructed. Algorithms 3 and 2 serve as a first step to the reconstruction of the graph corresponding to a stationary VAR(1) model, which even though is not the main subject of this paper, still could inspire future works.
@@ 4.1 Definition and properties
[…]
@@
1. There exists a directed path in $G$ between $i$ and $j$.
2. There exists $l\neq i,j\in V$ such that there are directed paths from $l$ to $i$ and $j$ respectively.
@@
Because the set of maximal classes is unique (Proposition 5), we can design Algorithm 1 to derive the set of maximal classes for any directed graph. This Algorithm first applies Kosaraju’s algorithm ([16]) to find all SCCs of the graph, and then construct the condensed graph, which is a directed acyclic graph where all SCCs are represented as a single node. Finally, we perform a Depth First Search (DFS) algorithm ([20]) on each of the nodes with in-degree zero in the condensed graph.
The computational complexity of Algorithm 1 is dictated by the Kosaraju’s algorithm, the construction of the condensed graph, and the subsequent DFS searches. Since these components are all linear operations relative to the size of the graph, the complexity of the algorithm is upper bounded by $O(k(n+E))$, where $n$ is the number of nodes, $E$ is the number of edges in the original graph, and $k$ is the number of SCCs in the condensed graph. Consequently, the best-case scenario achieves a complexity of $O(n+E)$, while the worst-case upper bound is $O(n(n+E))$.
Algorithm 1 Algorithm to identify maximal classes of a directed graph
@@
[…]
12:    $j\leftarrow j+1$   return j1j-1 (number of maximal classes) and maxc (list of maximal classes)
To reconstruct the list of graphs associated with a given set of maximal classes, that is, the inverse problem of the precedent algorithm, we propose Algorithm 2. This algorithm starts with identifying candidate sources for each maximal class. It then iterates through the complete set of possible edges $\left(u,v\right)\in V\times V$ to isolate ’forbidden’ edges — those whose inclusion would violate the definitions of the maximal classes or their sources. Finally, the algorithm evaluates subsets of the remaining ’allowable’ edges, applying Algorithm 1 to verify which resulting graphs yield the target set of maximal classes.
The complexity of Algorithm 2 is significantly higher due to its combinatorial nature. While the initial source allocation and edge pruning is efficient at $O(m\cdot n^{2})$ with $m$ being the number of maximal classes, the subsequent exploration of edge subsets results in a worst-case complexity of $O(2^{p}\cdot n(n+E))$, where $p$ is the number of ’allowable’ edges. The resulting complexity is upper bounded by $O(m\cdot n^{2}+2^{p}\cdot n(n+E))$.
An R implementation of both the algorithms is available at https://github.com/Bi-xuan/maximal_class. While the procedures presented in this section provide a systematic way to reconstruct maximal classes from their corresponding graph structures as well as the inverse, it is important to note that these implementations are not computationally optimal and leave significant room for algorithmic refinement. However, such optimizations and a systematic empirical validation across a broad selection of synthetic data and large-scale simulated ecological networks remain outside the primary theoretical focus of this work. These enhancements and performance testing are dedicated to future research.
Algorithm 2 Algorithm for the reconstruction of graphs via maximal classes
Input: a list of maximal classes $\mathfrak{MC}=\left\{\mathcal{MC}_{k},k=1,\cdots,m\right\}$
   Output: the list of graphs associated with $\mathfrak{MC}$
1:procedure RECONSTRUCT($\mathfrak{MC}$)
2:  $V\leftarrow\bigcup_{k=1}^{m}\mathcal{MC}_{k}$ $\triangleright$ Define the set of nodes
3:  SourceCandidates $\leftarrow$ an empty list of lists
4:  for all $\mathcal{MC}_{k}\in\mathfrak{MC}$ do
5:    $S_{k}\leftarrow\left\{v\in\mathcal{MC}_{k}\mid v\notin\mathcal{MC}_{j},\forall j\neq k\right\}$ $\triangleright$ Identify nodes that belong to only one maximal class
6:    Add $S_{k}$ to SourceCandidates
7:  ForbiddenEdges $\leftarrow\emptyset$
8:  for all $(u,v)\in V\times V$ do
9:    for all $\mathcal{MC}_{k}\in\mathfrak{MC}$ do
10:     if $u\in\mathcal{MC}_{k}$ and $v\notin\mathcal{MC}_{k}$ then
11:      Add $(u,v)$ to ForbiddenEdges $\triangleright$ By definition of maximal classes
12:     if $v\in S_{k}$ and $u\notin S_{k}$ then
13:      Add $(u,v)$ to ForbiddenEdges $\triangleright$ Source nodes must have in-degree zero from outside of their SCC
14:  PossibleEdges $\leftarrow V\times V\backslashForbiddenEdges$
15:  PossibleGraphs $\leftarrow\emptyset$
16:  for all selection of edges $E\subseteq$ PossibleEdges do $\triangleright$ Permute over subsets of possible edges
17:    $G\leftarrow(V,E)$
18:    if MAXC($G$) $\mathfrak{MC}$ then $\triangleright$ Verify using Algorithm 1
19:     Add $G$ to PossibleGraphs       return PossibleGraphs
@@ 4.2 Derivation of maximal classes from $\Sigma$
[…]
@@
The proof is given in Appendix G. ∎
@@
$\mathfrak{MC}=\mathfrak{C}\backslash\mathfrak{C}^{\prime}$,
@@
$\mathfrak{C}$ $=\left\{\mathcal{C}_{i}\mid i\in[n]\right\},$
$\mathfrak{C}^{\prime}$ $=\left\{\mathcal{C}_{i}\in\mathfrak{C}\mid i\in[n],\mbox{ and }\exists k,l\in\mathcal{C}_{i}\ s.t.\ \sigma_{kl}=0\right\}.$
@@
1. There exists a directed path from $i$ to $j$, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction.
2. There exists a directed path from $j$ to $i$, then ${\operatorname{deg}}_{+}(i)\neq 0$, which means that the source of $\mathcal{MC}_{[i]}$ is a SCC containing $i$. Hence either $j$ belongs to the SCC, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction, or $j$ does not belong to the SCC, then the in-degree of the SCC is non-zero, i.e. the SCC containing $i$ is not the source of $\mathcal{MC}_{[i]}$, which is also a contradiction.
3. There exists another node $c\in[n]$ s.t. there are paths from $c$ to $i$, and from $c$ to $j$ respectively, then ${\operatorname{deg}}_{+}(i)\neq 0$, which means that the source of $\mathcal{MC}_{[i]}$ is a SCC containing $i$. Hence either $c$ belongs to the SCC, then $j\in\mathcal{MC}_{[i]}$, which is a contradiction, or $c$ does not belong to the SCC, then the in-degree of the SCC is non-zero, i.e. the SCC containing $i$ is not the source of $\mathcal{MC}_{[i]}$, which is also a contradiction.
@@
$\mathfrak{MC}\subseteq\mathfrak{C}\backslash\mathfrak{C}^{\prime}$.
Let $\mathcal{C}_{i}\in\mathfrak{C}\backslash\mathfrak{C}^{\prime}$ for certain $i\in[n]$ and assume that $\mathcal{C}_{i}\notin\mathfrak{MC}$ (proof in Appendix F). We know that $i$ must belong to one maximal class in $\mathfrak{MC}$. Denote the maximal class containing $i$ as $\mathcal{MC}_{i}$. Then by Proposition 7,
@@
$\mathfrak{MC}=\mathfrak{C}\backslash\mathfrak{C}^{\prime}$.
@@
Algorithm 3 Algorithm for the reconstruction of maximal classes via the support of $\Sigma$
@@
12:     $\mathfrak{MC}\leftarrow\mathfrak{MC}\backslash\mathfrak{MC}[i]$ $\triangleright$ Keep only one of the repetitive elements
@@
$\mathfrak{C}=\left\{\{1,2\},\{1,2,3\},\{2,3\}\right\},$
@@
$\mathfrak{C}^{\prime}=\left\{\{1,2,3\}\right\},$
@@ 5.1 Maximal class characterization
$\psi_{G}\left(B\right)^{\left(\lambda_{st}\right)}=$ $\left[\begin{array}[]{cccc}\begin{array}[]{ccccc}0&\cdots&\left(\Sigma\Lambda\right)_{s1}&\cdots&0\end{array}&\cdots&\begin{array}[]{ccccc}\left(\Sigma\Lambda\right)_{s1}&\cdots&\left(\Sigma\Lambda\right)_{st}&\cdots&\left(\Sigma\Lambda\right)_{sn}\end{array}$$\cdots\end{array}\right$.
@@
  • Case 1.
    without loss of generality, assume the path is from $s$ to $a$, then there exist directed paths: $s\rightsquigarrow t\rightsquigarrow j$, and $s\rightsquigarrow a\rightsquigarrow i$. It contradicts the assumption.
  • Case 2.
    assume there exist directed paths: $l\rightsquigarrow s$, and $l\rightsquigarrow a$, then there exist directed paths: $l\rightsquigarrow a\rightsquigarrow i$, and $l\rightsquigarrow s\rightarrow t\rightsquigarrow j$. It also contradicts the assumption.
Therefore, for all $a\in\left[n\right]$ and $k\in\mathbb{N}$, $\left(\Sigma\Lambda\right)_{sa}\lambda_{ai}^{\left[k\right]}\lambda_{tj}^{\left[k\right]}=0$. The second term of the sum is also 0 by similar arguments. Hence, for all $s,t\in\left[n\right]$ s.t. $\left(s,t\right)\in\mathfrak{E}_{G}$,
@@
The proof is given in Appendix H. ∎
@@
$n_{c}^{\prime}$ $=|\left\{\{a,b\}\mid a,b\in\left[n\right],a,b\mbox{ belong to the same maximal class}\right\}|.$
@@
Let $A,B\in\mathbb{R}^{n}$. The notation $A\perp\!\!\!\perp B$ indicates that $A$ and $B$ are linearly independent, and $A\not\!\perp\!\!\!\perp B$ indicates that $A$ and $B$ are not linearly independent. Then, there exist $a^{\prime},b^{\prime}\in[n]$ where $\{a^{\prime},b^{\prime}\}\neq\{a,b\}$ s.t.
@@
$\mathbf{J}_{G_{1}}^{\left[\sigma_{ab}\right]}\not\!\perp\!\!\!\perp\mathbf{J}_{G_{1}}^{\left[\sigma_{a^{\prime}b^{\prime}}\right]},$
@@
$\mathbf{J}_{G_{2}}^{\left[\sigma_{ab}\right]}\not\!\perp\!\!\!\perp\mathbf{J}_{G_{2}}^{\left[\sigma_{a^{\prime}b^{\prime}}\right]}.$
@@
1. There exist $i,j\in[n]$ s.t. $i,j$ belong to the same maximal class in $G_{k_{1}}$, but do not in $G_{k_{2}}$.
2. There exist $\in[n] s.t. s,t$s,t belong to the same maximal class in $G_{k_{2}}$, but do not in $G_{k_{1}}$.
@@ 5.2 The dimension of the model for graphs without multi-edges
${\operatorname{dim}}\left(\mathbf{M}_{G}\right)=\mbox{rank}\left(\mathbf{J}_{G}\right)=\min\left\{n_{r},n_{c}^{\prime}\right\}=10.$
The proof of Theorem 5 is technical, and aligns with generic settings throughout this paper. It is divided into two separate cases: $n_{r}\leq n_{c}^{\prime}$ and $n_{r}>n_{c}^{\prime}$. We only introduce the proof for the first case, since the proof of the second case resembles the first one. For a complete proof of the second case, see Appendix J.
@@
[…]
@@
where $\left\{k,l\right\}\in\mathfrak{C}_{G}^{mc}$. Note that $\mathfrak{C}_{G}^{mc}\neq\emptyset$ by Lemma 8. In fact, this submatrix is a block matrix after reordering rows and columns. Apply the the Guttman rank additivity formula (see [8]), this matrix is full rank if and only if $B_{G}$ is full rank, up to a multiplication of a diagonal matrix with non-zero diagonals. For complete proof, see Appendix I. ∎
@@
1. $\left(k,a\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$.
2. $\left(a,k\right),\left(a,l\right)\in\mathfrak{E}_{G}^{\prime}$ and $\left(k,l\right),\left(l,k\right)\notin\mathfrak{E}_{G}^{\prime}$.
@@
is rational. From Appendix B, we know that the parametrization map
@@
Case 1.
Consider $\Lambda_{0}=\left(\lambda_{ij}^{0}\right)$ that satisfies:
@@
Case 2.
Let $G_{0}=\left(V,\mathfrak{E}_{G_{0}}\right)$ where
@@
Proof of case 11 in Theorem 5.
This is a direct result of Lemmas 9 and $1$. ∎